OpenStudy (anonymous):

A teacher has determined that the number of students in her Trig class has increased over a 5-year period at a rate of r given by the equation 5r²+14r-3=0. find r We got r=2 is this right?

OpenStudy (harsha19111999):

no

OpenStudy (harsha19111999):

you get 1/5

OpenStudy (anonymous):

harsha can u kindly show how u got the answer?

OpenStudy (anonymous):

we had to use quadratic formula.

OpenStudy (harsha19111999):

You have to factorize

OpenStudy (harsha19111999):

Hold on

OpenStudy (harsha19111999):

the equation should be written as 5r^2+15r-r-3

OpenStudy (harsha19111999):

now take 5r and -1 common. then u get 5r(r+3)-1(r+3) then r=-3 or r=1/5 r cant be -ve value

OpenStudy (anonymous):

using quadratic formula the equation that we had \[\frac{ -14+or-\sqrt{14^{2}-4(5)(-3)} }{ 10 }\]

OpenStudy (anonymous):

i get what harsha is saying but zuru i have no idea what u just said.

OpenStudy (anonymous):

i just know it can not be a portion of a student, so 1/5 cant be right

OpenStudy (anonymous):

so cant be -3 then u say both answers are wrong?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so there is no correct answer for this question?

OpenStudy (anonymous):

there is, it just has to be done using quadratic formula. and I am not sure I got it right

OpenStudy (anonymous):

should I ask again using the formula? Im new here so not sure what to do

OpenStudy (anonymous):

yes please